Part A. The maximum contaminant level (MCL) of chlorobenzene (C6H5Cl) is 0.1 mg/L in water. Express this MCL in terms of (a) ppm, (b) ppb, and (c) moles/m3.

Part B. The maximum contaminant level (MCL) of benzene (C6H6) is 0.005 mg/L in water. Express this MCL in terms of (a) ppm, (b) ppb, and (c) moles/m3.

This is one question. Thank you.
Answer #1

1 ppm = 1 mg/L

1 ppb = 10-3 mg/L

Part A:

0.1 mg/L of chlorobenzene = 0.1 ppm of chlorobenzene

0.1 mg/L of chlorobenzene = 0.1 x 1000 = 100 ppb of chlorobenzene

0.1 mg/L = 0.1 x 10-3 g/L = 0.1 x 10-3 g / 10-3 m3 = 0.1 g/m3

Molar mass of chlorobenzene (C6H5Cl) = 112.56 g/mol

0.1 g of chlorobenzene = 0.1/112.56 = 8.88 x 10-4 mol

Hence, 0.1 mg/L = 0.1 g/m3 = 8.88 x 10-4 mol/m3

-------------------------

Part B:

0.005 mg/L of benzene = 0.005 ppm of benzene

0.1 mg/L of benzene = 0.005 x 1000 = 5 ppb of benzene

0.005 mg/L = 0.005 x 10-3 g/L = 0.005 x 10-3 g / 10-3 m3 = 0.005 g/m3

Molar mass of benzene = 78.11 g/mol

0.005 g of benzene (C6H6) = 0.005/78.11 = 6.40 x 10-5 mol

Hence, 0.005 mg/L = 0.005 g/m3 = 6.40 x 10-5 mol/m3